I try to challenge myself and my friends from time to time but i seem to need some help on these problems.
1.(3s+3)squared
2.(5z+1)squared
3.(10y+1)squared
4.(1+10y)squared

Answer by stanbon(75887)

1.(3s+3)squared
2.(5z+1)squared
3.(10y+1)squared
4.(1+10y)squared
-------------
Here's the form for squaring a binomial:
(a+b)^2 = a^2 + 2ab + b^2
----
Your 1st problem:
a = 2s, b = 3
Ans: (2s)^2 + 2(2s)(3) + 3^2 = 4s^2 = 12s + 9
---
Do the same for 2 thru 4.
Cheers,
Stan H.